OpenStudy (anonymous):

Members of a school social service group hold a spaghetti dinner to raise money for orphan children. They must sell tickets at a price that will earn them more money than the cost of food. The formula for this scenario is t = n (p − c) where t = total profit earned from the event c = cost of food per plate n = number of tickets sold p = price charged for each dinner Rewrite the equation to determine the number of tickets they need to sell.

OpenStudy (anonymous):

to rewrite the equation in a way to determin the number of tickets you need to get n by itself

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

can u work through the problem with me???

OpenStudy (anonymous):

if i dont get this right ill be grounded for like moth maybe even summer

OpenStudy (anonymous):

im not going to give you the answer but yes i will work through it when it says to rewrite it to determine the number of tickets ...it means it wants n by itself and everything else on the other side. bc n represents the number of tickets sold the equation is t=n (p-c) n is being multiplied by (p-c) ..to get n by itself all you need to do is make (p-c) on the other side

OpenStudy (anonymous):

A. n x t = (p-c) B. n= t/p-c C. t/n=(p-c) D. p-c/t=1/n

OpenStudy (anonymous):

so how do you cancel out multplication when doing equations?

OpenStudy (anonymous):

i dont remeber that y i need help

OpenStudy (anonymous):

you divide what you need to disappear. but when you divide it on one side it has to be divided on the other side as well

OpenStudy (anonymous):

i still dont know im sorry im not verry smart

OpenStudy (anonymous):

 notice its divided on both sides?